OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points (2, 5) and (8, 3)? (1 point)The fraction one third The fraction negative one third 3 –3 please help im new to this and I will fan medal and greatly appreciate :)

OpenStudy (hockeychick23):

y2-y1/x2-x1 is the formula you need to use, so it should look like this: 3-5/8-2= -2/6= -1/3

OpenStudy (anonymous):

The formula to find the slope of a linear line is (y2 - y1)/(x2 - x1) So plug in your values... (5 - 3)/(2 - 8) = 2/-6 Simplify... 2/-6 = 1/-3 (rise/run) And there you go (: Easy peasy.